OpenStudy (anonymous):

f(x)= 2 * 5^x which facts are true about the functions a. It is increasing b.The y intercept is (0,5) c.The y intercept is (0,2) d.The domain of f(x) is all real numbers e.The range of f(x) is y>5 f.It is decreasing

OpenStudy (anonymous):

\[f(x)=2\times 5^x\]?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

definitely increasing because it is an exponential function and the base(5) is bigger than zero

OpenStudy (anonymous):

so a is correct.

OpenStudy (anonymous):

b is wrong because \[f(0)=2\times 5^0=2\]

OpenStudy (anonymous):

the above shows that c is correct as well.

OpenStudy (anonymous):

d is also correct. you can raise 5 to any power that you wish

OpenStudy (anonymous):

e is incorrect. we already know that \[f(0)=2\] so it makes no sense to say the range is \[y>5\]

OpenStudy (anonymous):

and if is incorrect because if it is increasing it cannot be decreasing!